# High-precision measurement of the atomic mass of 84Sr and implications to isotope shift studies

### v1

@@ IV Conclusion and outlook
In summary, we have performed direct high-precision mass and double-beta decay value measurements of 84Sr using the PI-ICR technique with the JYFLTRAP double PTMS. The mass and values were determined with a precision approximately 30 times higher than those adopted in AME2020, aligning well with the literature. The new mass value was utilized to conduct an updated King plot analysis for transitions at 689 nm and 698 nm for 84,86,87Sr in relation to 88Sr. Our results showcase an enhanced nonlinear King plot, in good agreement with previously reported values, offering smaller uncertainties. The current contribution of the mass uncertainty to the King plot analysis enables future precision enhancements by several orders of magnitude in optical spectroscopy, mitigating statistical and systematic errors in both transitions. This potential advancement could lead to unparalleled sensitivity to new physics. More precise isotope shift measurements of Sr isotopes, providing stringent experimental constraints on King linearity, are highly required. Performing such a measurement with a radioactive isotope e.g., 90Sr, to avoid complications due to hyperfine structure, is feasible at the IGISOL facility. In conjunction with additional measurements of isotope shifts in elements such as calcium, this will contribute to enhancing precision in atomic structure calculations and imposing constraints on new physics.
Acknowledgements.
@@
We acknowledge the staff of the Accelerator Laboratory of University of Jyväskylä (JYFL-ACCLAB) for providing stable online beam. We thank the support by the Academy of Finland under the Finnish Centre of Excellence Programme 2012-2017 (Nuclear and Accelerator Based Physics Research at JYFL) and projects No. 306980, No. 312544, No. 275389, No. 284516, No. 295207, No. 314733, No. 315179, No. 327629, No. 320062, No. 354589 and 345869. The support by the EU Horizon 2020 research and innovation program under grant No. 771036 (ERC CoG MAIDEN) is acknowledged. We express gratitude for the productive results stemming from conversations with R. de Groote and X. F. Yang.References
- King (1984) W. H. King, Isotope Shifts in Atomic Spectra (Plenum, New York, 1984).
- Flambaum et al. (2018) V. V. Flambaum, A. J. Geddes, and A. V. Viatkina, Phys. Rev. A 97, 032510 (2018).
- Dammalapati et al. (2009) U. Dammalapati, S. De, K. Jungmann, et al., Eur. Phys. J. D 53, 1 (2009).
- Shi et al. (2017) C. Shi, F. Gebert, C. Gorges, et al., Appl. Phys. B 123, 2 (2017).
- Nazé et al. (2015) C. Nazé, J. G. Li, and M. Godefroid, Phys. Rev. A 91, 032511 (2015).
- Frugiuele et al. (2017) C. Frugiuele, E. Fuchs, G. Perez, and M. Schlaffer, Phys. Rev. D 96, 015011 (2017).
- Berengut et al. (2018a) J. C. Berengut, D. Budker, C. Delaunay, V. V. Flambaum, C. Frugiuele, E. Fuchs, C. Grojean, R. Harnik, R. Ozeri, G. Perez, and Y. Soreq, Phys. Rev. Lett. 120, 091801 (2018a).
- Miyake et al. (2019) H. Miyake, N. C. Pisenti, P. K. Elgee, et al., Phys. Rev. Research 1, 033113 (2019).
- Manovitz et al. (2019) T. Manovitz, R. Shaniv, Y. Shapira, R. Ozeri, and N. Akerman, Phys. Rev. Lett. 123, 203001 (2019).
- Knollmann et al. (2019) F. W. Knollmann, A. N. Patel, and S. C. Doret, Phys. Rev. A 100, 022514 (2019).
- Counts et al. (2020a) I. Counts, J. Hur, D. P. L. Aude Craik, H. Jeon, C. Leung, J. C. Berengut, A. Geddes, A. Kawasaki, W. Jhe, and V. Vuletić, Phys. Rev. Lett. 125, 123002 (2020a).
- Solaro et al. (2020) C. Solaro, S. Meyer, K. Fisher, J. C. Berengut, E. Fuchs, and M. Drewsen, Phys. Rev. Lett. 125, 123003 (2020).
- Nesterenko et al. (2020) D. Nesterenko, R. de Groote, T. Eronen, Z. Ge, M. Hukkanen, A. Jokinen, and A. Kankainen, International Journal of Mass Spectrometry 458, 116435 (2020).
- Counts et al. (2020b) I. Counts, J. Hur, D. P. L. Aude Craik, H. Jeon, C. Leung, J. C. Berengut, A. Geddes, A. Kawasaki, W. Jhe, and V. Vuletić, Phys. Rev. Lett. 125, 123002 (2020b).
- Rehbehn et al. (2021) N.-H. Rehbehn, M. K. Rosner, H. Bekker, et al., Phys. Rev. A 103, L040801 (2021).
- Müller et al. (2021) R. A. Müller, V. A. Yerokhin, A. N. Artemyev, and A. Surzhykov, Phys. Rev. A 104, L020802 (2021).
- Hur et al. (2022) J. Hur, D. P. L. A. Craik, I. Counts, et al., Phys. Rev. Lett. 128, 163201 (2022).
- Figueroa et al. (2022) N. L. Figueroa, J. C. Berengut, V. A. Dzuba, et al., Phys. Rev. Lett. 128, 073001 (2022).
- Ono et al. (2022) K. Ono, Y. Saito, T. Ishiyama, et al., Phys. Rev. X 12, 021003 (2022).
- Stellmer et al. (2014) S. Stellmer, F. Schreck, and T. C. Killian, in Annual Review of Cold Atoms and Molecules, Vol. 2, edited by K. W. Madison, K. Bongs, L. D. Carr, A. M. Rey, and H. Zhai (World Scientific, Singapore, 2014) Chap. 1.
- Rana et al. (2012) R. Rana, M. Höcker, and E. G. Myers, Phys. Rev. A 86, 050502 (2012).
- Wang et al. (2021) M. Wang, W. Huang, F. Kondev, G. Audi, and S. Naimi, Chinese Physics C 45, 030003 (2021).
- Kondev et al. (2021) F. Kondev, M. Wang, W. Huang, S. Naimi, and G. Audi, Chinese Physics C 45, 030001 (2021).
- Eronen et al. (2012) T. Eronen, V. S. Kolhinen, V. V. Elomaa, D. Gorelov, U. Hager, J. Hakala, A. Jokinen, A. Kankainen, P. Karvonen, S. Kopecky, I. D. Moore, H. Penttilä, S. Rahaman, S. Rinta-Antila, J. Rissanen, A. Saastamoinen, J. Szerypo, C. Weber, and J. Äystö, European Physical Journal A 48, 1 (2012).
- Eronen et al. (2016) T. Eronen, A. Kankainen, and J. Äystö, Progress in Particle and Nuclear Physics 91, 259 (2016).
- Moore et al. (2013) I. D. Moore, T. Eronen, D. Gorelov, J. Hakala, A. Jokinen, A. Kankainen, V. S. Kolhinen, J. Koponen, H. Penttilä, I. Pohjalainen, M. Reponen, J. Rissanen, A. Saastamoinen, S. Rinta-Antila, V. Sonnenschein, and J. Äystö, Nuclear Instruments and Methods in Physics Research, Section B: Beam Interactions with Materials and Atoms 317, 208 (2013).
- Karvonen et al. (2008) P. Karvonen, I. D. Moore, T. Sonoda, T. Kessler, H. Penttilä, K. Peräjärvi, P. Ronkanen, and J. Äystö, Nuclear Instruments and Methods in Physics Research, Section B: Beam Interactions with Materials and Atoms 266, 4794 (2008).
- Nieminen et al. (2001) A. Nieminen, J. Huikari, A. Jokinen, J. Äystö, P. Campbell, and E. C. Cochrane, Nuclear Instruments and Methods in Physics Research, Section A: Accelerators, Spectrometers, Detectors and Associated Equipment 469, 244 (2001).
- Savard et al. (1991) G. Savard, S. Becker, G. Bollen, H. J. Kluge, R. B. Moore, T. Otto, L. Schweikhard, H. Stolzenberg, and U. Wiess, Physics Letters A 158, 247 (1991).
- Nesterenko et al. (2018) D. A. Nesterenko, T. Eronen, A. Kankainen, L. Canete, A. Jokinen, I. D. Moore, H. Penttilä, S. Rinta-Antila, A. de Roubin, and M. Vilen, European Physical Journal A 54, 0 (2018).
- Nesterenko et al. (2021) D. A. Nesterenko, T. Eronen, Z. Ge, A. Kankainen, and M. Vilen, Eur. Phys. J. A 57, 302 (2021).
- Eliseev et al. (2014) S. Eliseev, K. Blaum, M. Block, A. Dörr, C. Droese, T. Eronen, M. Goncharov, M. Höcker, J. Ketter, E. M. Ramirez, D. A. Nesterenko, Y. N. Novikov, and L. Schweikhard, Applied Physics B: Lasers and Optics 114, 107 (2014).
- Eliseev et al. (2013) S. Eliseev, K. Blaum, M. Block, C. Droese, M. Goncharov, E. Minaya Ramirez, D. A. Nesterenko, Y. N. Novikov, and L. Schweikhard, Physical Review Letters 110, 82501 (2013).
- Kramida et al. (2020) A. Kramida, Yu. Ralchenko, J. Reader, and and NIST ASD Team, NIST Atomic Spectra Database (ver. 5.8), [Online]. Available: https://physics.nist.gov/asd [2021, January 19]. National Institute of Standards and Technology, Gaithersburg, MD. (2020).
- Kellerbauer et al. (2003) A. Kellerbauer, K. Blaum, G. Bollen, F. Herfurth, H. J. Kluge, M. Kuckein, E. Sauvan, C. Scheidenberger, and L. Schweikhard, European Physical Journal D 22, 53 (2003).
- Roux et al. (2013) C. Roux, K. Blaum, M. Block, C. Droese, S. Eliseev, M. Goncharov, F. Herfurth, E. M. Ramirez, D. A. Nesterenko, Y. N. Novikov, and L. Schweikhard, The European Physical Journal D 67, 1 (2013).
- Birge (1932) R. T. Birge, Physical Review 40, 207 (1932).
- Huang et al. (2021) W. Huang, M. Wang, F. Kondev, G. Audi, and S. Naimi, Chinese Physics C 45, 030002 (2021).
- Kellerbauer et al. (2007) A. Kellerbauer, G. Audi, D. Beck, K. Blaum, G. Bollen, C. Guénaut, F. Herfurth, A. Herlert, H.-J. Kluge, D. Lunney, S. Schwarz, L. Schweikhard, C. Weber, and C. Yazidjian, Phys. Rev. C 76, 045504 (2007).
- Haettner et al. (2011) E. Haettner, D. Ackermann, G. Audi, K. Blaum, M. Block, S. Eliseev, T. Fleckenstein, F. Herfurth, F. P. Heßberger, S. Hofmann, J. Ketelaer, J. Ketter, H.-J. Kluge, G. Marx, M. Mazzocco, Y. N. Novikov, W. R. Plaß, S. Rahaman, T. Rauscher, D. Rodríguez, H. Schatz, C. Scheidenberger, L. Schweikhard, B. Sun, P. G. Thirolf, G. Vorobjev, M. Wang, and C. Weber, Phys. Rev. Lett. 106, 122501 (2011).
- Booij et al. (1971) H. Booij, E. Van Hoek, H. Van Der Molen, W. Slot, and J. Blok, Nuclear Physics A 160, 337 (1971).
- Morton et al. (1971) J. Morton, W. Davies, W. Mclatchie, D. W., and J. Kitching, Nuclear Physics A 161, 228 (1971).
- Takano et al. (2017) T. Takano, R. Mizushima, and H. Katori, Applied Physics Express 10, 072801 (2017).
- King (1963) W. H. King, J. Opt. Soc. Am. 53, 638 (1963).
- Berengut et al. (2018b) J. C. Berengut, D. Budker, C. Delaunay, V. V. Flambaum, C. Frugiuele, E. Fuchs, C. Grojean, R. Harnik, R. Ozeri, G. Perez, and Y. Soreq, Phys. Rev. Lett. 120, 091801 (2018b).
- York et al. (2004) D. York, N. M. Evensen, M. L. Martınez, and J. De Basabe Delgado, American Journal of Physics 72, 367 (2004).
- Ludlow et al. (2015) A. D. Ludlow, M. M. Boyd, J. Ye, E. Peik, and P. O. Schmidt, Rev. Mod. Phys. 87, 637 (2015).

### v2

@@ IV Conclusion and outlook
In summary, we have performed direct high-precision mass and double-beta decay value measurements of 84Sr using the PI-ICR technique with the JYFLTRAP double PTMS. The newly refined mass and values for 84Sr were in agreement with the adopted values in AME2020, while a precision approximately 30 times higher than those adopted in AME2020 were achieved. The new mass value was utilized to conduct an updated King plot analysis for transitions at 689 nm and 698 nm for 84,86,87Sr in relation to 88Sr. Our results reveal a nonlinear King plot, consistent with prior research and featuring reduced uncertainties. The current contribution of the mass uncertainty to the King plot analysis enables future precision enhancements by several orders of magnitude in optical spectroscopy, mitigating statistical and systematic errors in both transitions. This potential advancement in both transitions could lead to unparalleled sensitivity to new physics. More precise isotope shift measurements of Sr isotopes, providing stringent experimental constraints on King linearity, are highly required. Performing such a measurement with a radioactive isotope e.g., 90Sr, to avoid complications due to hyperfine structure, is feasible at the IGISOL facility. In conjunction with additional measurements of isotope shifts in elements such as calcium, this will contribute to testing the predictions from atomic theory and imposing constraints on new physics.
V Acknowledgements
@@
We acknowledge the staff of the Accelerator Laboratory of University of Jyväskylä (JYFL-ACCLAB) for providing stable online beam. We thank the support by the Academy of Finland under the Finnish Centre of Excellence Programme 2012-2017 (Nuclear and Accelerator Based Physics Research at JYFL) and projects No. 306980, No. 312544, No. 275389, No. 284516, No. 295207, No. 314733, No. 315179, No. 327629, No. 320062, No. 354589 and 345869. The support by the EU Horizon 2020 research and innovation program under grant No. 771036 (ERC CoG MAIDEN) is acknowledged. We express gratitude for the productive results stemming from conversations with R. de Groote and X. F. Yang.